OpenStudy (anonymous):

Need help solving simple equations 3x-2 x^2-4x+1

OpenStudy (anonymous):

Well you have to set each one equal to 0

OpenStudy (anonymous):

And solve for x

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

So for the first one, it is 3x-2=0. Then you add 2 to the other side. Then you divide by 3. That gives you x=2/3

OpenStudy (anonymous):

For the second one, you have to factor it

OpenStudy (anonymous):

wait

OpenStudy (freckles):

I don't see any equations above.

OpenStudy (freckles):

We can't solve an equation without an equation.

OpenStudy (freckles):

In fact, the things you have written look like two separate expressions.

OpenStudy (anonymous):

For the second one, I think you need the quadratic formula

OpenStudy (freckles):

@tookie298 What are the instructions?

OpenStudy (anonymous):

if f(x)=3x-2 and g(x)=x^2-4x+1,find (f+g)(x)

OpenStudy (freckles):

Ok then you add f and g (3x-2)+(x^2-4x+1)

OpenStudy (freckles):

then combine like terms

OpenStudy (anonymous):

thats it?

OpenStudy (freckles):

that is it

OpenStudy (anonymous):

someone told me I needed to solve them individually than combine together

OpenStudy (freckles):

if you are given f and g, and the only thing you are asked to do is find f+g then yeah that is it that is it being all you have to do is f+g (3x-2)+(x^2-4x+1)

OpenStudy (anonymous):

But if I wanted to solve it individually I would set 3x-2 to 3x-2=0?

OpenStudy (anonymous):

I got the answer to the question but I'm not sure how that answer is correct

OpenStudy (freckles):

I don't know what you when you say solve 3x-2 The sentence "Solve 3x-2." doesn't make any sense because there is nothing to solve. 3x-2 is not an equation. You can not solve it.

OpenStudy (freckles):

If you were given 3x-2=0 <--this is an equation which can be solved

OpenStudy (anonymous):

this is a expression than

OpenStudy (anonymous):

the answer is x^2-x-1 but how did they come up with x^2?

OpenStudy (freckles):

again you are suppose to add f and g

OpenStudy (freckles):

(3x-2)+(x^2-4x+1)

OpenStudy (freckles):

i see an x^2

OpenStudy (freckles):

re order the expression 3x-2+x^2-4x+1 x^2+3x-4x+1 can you finish the problem now?

OpenStudy (anonymous):

Yeah i get that

OpenStudy (triciaal):

you cannot have an equation without an equal sign equation means the 2 expressions are equivalent

OpenStudy (freckles):

Oops forgot to bring down the -2

OpenStudy (freckles):

x^2+3x-4x+1-2

OpenStudy (anonymous):

Thanks @triciaal

OpenStudy (anonymous):

I understand that freckles

OpenStudy (freckles):

so you see how they get the answer then?

OpenStudy (anonymous):

I see how they got but I don't understand how to come to that answer myself

OpenStudy (anonymous):

got it*

OpenStudy (anonymous):

I mean i understand x^2 but I don't see how it went from "+" to "-"

OpenStudy (freckles):

Ok so do you know what + means? + means to add You are asked to find f+g f=3x-2 g=x^2-4x+1 replace f with 3x-2 and replace g with x^2-4x+1 f+g (3x-2)+(x^2-4x+1) Remember + just we need to add the expreesions

OpenStudy (freckles):

Reorder the terms so that like terms are together x^2+3x-4x-2+1

OpenStudy (anonymous):

I apologize for my lack of knowledge pertaining to this its been awhile

OpenStudy (freckles):

what is 3-4?

OpenStudy (anonymous):

I know its -x is that how the expression went from x^2"+"3x-4x-2+1 too x^2(-)x+1? Thats where I've been confused at

OpenStudy (freckles):

it doesn't it goes from x^2+3x-4x-2+1 to x^2-x-1

OpenStudy (anonymous):

yeah thats what I ment the "+">"-" part

OpenStudy (anonymous):

the addition to subtract part

OpenStudy (freckles):

well 3x-4x=-1x and -2+1=-1

OpenStudy (anonymous):

Ah...I thought so

OpenStudy (anonymous):

Thanks @freckles

OpenStudy (triciaal):

you are just grouping like terms and simplifying all the terms with x^2 plus all the terms in x plus all the constants

OpenStudy (freckles):

lol I'm still confused were you were confused at